1.What is the magnitude and direction of the field at 0.25 m from a -5.6×10^-6 C point charge.



1
Expert's answer
2022-02-04T07:42:04-0500

The field is directed toward the charge, the magnitude is


"|E|=\dfrac1{4\pi\epsilon_0}\dfrac q{r^2}=805285\text{ V/m}."